### v1

# Invariant Measures in Time-Delay Coordinates for Unique Dynamical System Identification

Abstract
Invariant measures are widely used to compare chaotic dynamical systems, as they offer robustness to noisy data, uncertain initial conditions, and irregular sampling. However, large classes of systems with distinct transient dynamics can still exhibit the same asymptotic statistical behavior, which poses challenges when invariant measures alone are used to perform system identification. Motivated by Takens’ seminal embedding theory, we propose studying invariant measures in time-delay coordinates, which exhibit enhanced sensitivity to the underlying dynamics. Our first result demonstrates that a single invariant measure in time-delay coordinates can be used to perform system identification up to a topological conjugacy. This result already surpasses the capabilities of invariant measures in the original state coordinate. Continuing to explore the power of delay-coordinates, we eliminate all ambiguity from the conjugacy relation by showing that unique system identification can be achieved using additional invariant measures in time-delay coordinates constructed from different observables. Our findings improve the effectiveness of invariant measures in system identification and broaden the scope of measure-theoretic approaches to modeling dynamical systems.
1 Introduction
DISTRIBUTION STATEMENT A. Approved for public release. Distribution is unlimited.
Coordinate transformations play essential roles across numerous fields of physics and applied mathematics, including planetary science [20], relativity [17], data-driven dynamical systems [12], and partial differential equations [22, 42]. Experimental data from complex physical processes is typically analyzed as time-trajectory data within certain “state” coordinate axes. In this paper, we leverage tools from ergodic theory [16] and delay embedding theory [53, 60] to show that one can equivalently view this trajectory data as a stationary point-cloud, i.e., a discrete probability measure, within a coordinate frame whose axes instead encode the system’s temporal evolution. This new perspective motivates the use of measure transport, a class of approaches for comparing probability measures which has revolutionized many areas of data science and machine learning, for analyzing experimental time-series data [4, 24, 47, 54].
In this work, we consider the problem of constructing data-driven models of dynamical systems, a fundamental challenge in applications such as gravitational wave parameter estimation [15], fluid flow surrogate modeling [32], ion thruster model calibration [25], and weather prediction [7]. Such models can facilitate insights into the underlying physics of complex systems and serve as valuable tools for performing state prediction and control [40]. The task of system identification is typically formulated as an optimization procedure, where empirical time-series data is compared to outputs of candidate models, often expressed as differential equations [1, 5, 6, 9, 38]. Sparse sampling, noisy measurements, partial observations, and chaotic dynamics represent common challenges in system identification, which can hinder one’s ability to discover physically informative models.
Directly comparing long time-series data from chaotic systems is known to be impractical, as the intrinsic unpredictability of chaos makes it difficult to differentiate between model inaccuracies and inherent dynamical instability [8, 29, 39, 65]. In contrast, the invariant measures of chaotic systems, which are independent of initial conditions and can exhibit smooth parameter dependence, offer a robust alternative [51, 67]. These measures are resilient to measurement noise and can be accurately approximated even under slow sampling conditions [8], making them valuable tools for identifying dynamical systems.
In recent years, several works have used invariant measures in data-driven dynamical system modeling. For instance, the Wasserstein distance, derived from optimal transport theory, has been used to compare dynamical attracrors [25, 41, 65]. Moreover, the parameter estimation inverse problem has been reformulated as a PDE-constrained optimization, utilizing the stationary solution of the Fokker-Planck equation as a surrogate for the invariant measure [8, 14, 65]. Pseudo-metrics based on harmonic time averages of observables were introduced for system comparison in [37]. Invariant measures, along with other quantities related to the long-term statistics, have been incorporated as regularization terms when training models according to a trajectory-based mean-squared error loss [29, 33, 48, 55, 62]. Other works have explored optimal perturbations to generate desired linear responses in invariant measures [3, 21, 31]. Furthermore, invariant measures and linear response theory have been applied to study variations in climate systems [27, 50]. Various works have also explored model architectures and loss functions which yield accurate invariant measure reconstructions, without directly leveraging the statistical properties of trajectories during training [11, 26, 34, 45].
Figure 1: Illustration of the difference between state-coordinate invariant measures and delay coordinate invariant measures for the torus rotation $T_{\alpha,\beta}(z_{1},z_{2})=(z_{1}+\alpha,z_{2}+\beta)\pmod{1}.$ The top row shows four different choices of $(\alpha,\beta)$ for which the state-coordinate invariant measures are exactly the same. However, all four systems can be distinguished by their invariant measure in time-delay coordinates, as shown in the bottom row.
Despite the advantages of using invariant measures to perform model identification, they face a fundamental limitation: systems with different transient dynamics may still exhibit identical long-term statistical behavior, as described by the same invariant measure [19]. Thus, invariant measures alone cannot fully distinguish between such systems. Motivated by Takens’ time-delay embedding theory [60] and its later development [53], we instead propose using invariant measures in time-delay coordinates for dynamical system identification. As illustrated in Figure 1, systems that share the same invariant measure in their original state-coordinates can display distinct measures in time-delay coordinates, offering greater insight into their underlying dynamics.
In this work, we present two key results. First, we prove that if two dynamical systems share the same invariant measure in time-delay coordinates, they are topologically conjugate on the support of their invariant measures. Second, we show that the non-uniqueness from the conjugacy relation can be eliminated by using multiple delay-coordinate invariant measures derived from different observables. That is, a finite set of delay-coordinate invariant measures provides enough information to uniquely identify a dynamical system under mild assumptions. A summary of these results is depicted in the flowchart in Figure 2. While prior studies have explored the numerical use of delay-coordinate invariant measures [8, 25, 39], our work is the first to provide theoretical guarantees supporting their use for system identification.
Invariant measure $T\#\mu=\mu$ One delay-coordinate invariant measure $\hat{\mu}_{(y,T)}^{(m+1)}=\Psi_{(y,T)}^{(m+1)}\#\mu$ $m$ delay-coordinate invariant measures $\big{\{}\hat{\mu}_{(y_{i},T)}^{(m+1)}\big{\}}_{i=1}^{m}$ No uniqueness Conjugacy on $\text{supp}(\mu)$ Uniqueness on $\text{supp}(\mu)$ (new coordinate)(more measures)(Theorem 3.1)(Theorem 3.2)+ initialcondition
Figure 2: Flowchart of our main results. While the invariant measure $\mu$ cannot uniquely identify the dynamical system, one delay-coordinate invariant measure can ensure topological conjugacy of the reconstructed system on $\text{supp}(\mu)$. Moreover, several delay-coordinate invariant measures can uniquely determine the dynamics on $\text{supp}(\mu)$, with an appropriate initial condition.
The paper is organized as follows: In Section 2, we review the necessary background on invariant measures and time-delay embedding. In Section 3, we present and prove our main results, along with discussions. We conclude in Section 4.
2 Background and Preliminaries
This section reviews the essential background and preliminary results necessary for the statements and proofs of our main results in Section 3. In Section 2.1, we cover pushforward measures, invariant measures, and basins of attraction. In Section 2.2, we introduce the time-delay map and the corresponding invariant measure in time-delay coordinates, which we propose to use for system identification. Finally, in Section 2.3, we discuss generalizations of the Whitney and Takens embedding theorems, which ensure that the invariant measure in time-delay coordinates is well-defined.
2.1 Invariant Measures
Throughout, we will formulate many of our definitions and results in the abstract setting of Polish spaces. A Polish space $X$ is a separable completely metrizable topological space. We will denote by $\mathcal{P}(X)$ the space of probability measures on a Polish space $X$, which consists of all Borel measures that assign unit measure to $X$. Throughout our proofs, it will be important to reference the support of a given probability measure, which can intuitively be viewed as the region on which the measure is concentrated; see [46, Theorem 2.1].
Definition 2.1 (Support of a measure).
If $X$ is a Polish space and $\mu\in\mathcal{P}(X)$ is a Borel probability measure, then the support of $\mu$ is the unique closed set $\text{supp}(\mu)\subseteq X$ such that $\mu(\text{supp}(\mu))=1$, and for any other closed set $C\subseteq X$ satisfying $\mu(C)=1$, it holds that $\text{supp}(\mu)\subseteq C$.
We will frequently need to discuss how probability measures are transformed under the action of a measurable function. This concept is formalized with the notion of a pushforward measure.
Definition 2.2 (Pushforward measure).
Let $X$ and $Y$ be Polish spaces, let $f:X\to Y$ be Borel measurable, and let $\mu\in\mathcal{P}(X).$ Then, the pushforward of $\mu$ is the probability measure $f\#\mu\in\mathcal{P}(Y)$, defined by $(f\#\mu)(B):=\mu(f^{-1}(B))$, for all Borel sets $B\subseteq Y$.
We next define the notion of an invariant measure, which is a probability measure that remains unchanged under the pushforward of a given self-transformation $T:X\to X$. In the study of dynamical systems, application of the self-transformation $T$ to a state $x\in X$, i.e., $x\mapsto T(x)$, represents the evolution of time. The invariant measure provides a statistical description of the system’s asymptotic behavior.
Definition 2.3 (Invariant measure).
Let $X$ be a Polish space and let $T:X\to X$ be Borel measurable. The probability measure $\mu\in\mathcal{P}(X)$ is said to be $T$-invariant (or simply invariant when the map $T$ is clear from context) if $T\#\mu=\mu$.
Additionally, the measure $\mu$ is said to be $T$-ergodic if $T^{-1}(B)=B$ implies that $\mu(B)\in\{0,1\}$, i.e., when all invariant sets trivially have either 00 or full measure. Given a dynamical system $T:X\to X$, it is natural to ask which initial conditions $x\in X$ produce trajectories $\{T^{k}(x)\}_{k\in\mathbb{N}}$ that yield empirical measures which are asymptotic to a given invariant measure. Given that experimentalists primarily observe trajectories of dynamical systems, it is important to understand the extent to which this data provides information about the underlying invariant measure. Towards this, we now define the notion of a basin of attraction.
Definition 2.4 (Basin of attraction).
Let $X$ be a Polish space, and let $T:X\to X$ be Borel measurable. The basin of attraction of a $T$-invariant measure $\mu\in\mathcal{P}(X)$ with compact support is
[…]
If the space $X$ admits a Lebesgue measure, then $\mu$ is said to be a physical measure when the basin of attraction (1) has positive Lebesgue measure. From the perspective of experimentalists, physical measures are the relevant invariant measures which can be observed during data collection. Under certain assumptions on the space $X$ and the system $T$, the existence of physical measures is known (see [67, Theorem 1]), with common examples including attracting fixed points, stable limit cycles, and various chaotic attractors, e.g., the Lorenz-63 system [35].
2.2 Invariant Measures in Time-Delay Coordinates
In this work, we compare two dynamical systems, $T:X\to X$ and $S:Y\to Y$. A fundamental notion of equivalence between the systems $T$ and $S$ is given by the concept of topological conjugacy. It defines an equivalence relation, preserving crucial properties of the dynamics, such as the structure of periodic orbits, topological entropy, and ergodicity [30].
Definition 2.5 (Topological conjugacy).
Let $X$ and $Y$ be Polish spaces, and let $T:X\to X$ and $S:Y\to Y$ be continuous. The maps $T$ and $S$ are said to be topologically conjugate if there exists a homeomorphism $h:X\to Y$, such that $S=h\circ T\circ h^{-1}$.
In Definition 2.5, the homeomorphism $h:X\to Y$ is known as the conjugating map. When $T:X\to X$ admits $\mu\in\mathcal{P}(X)$ as an invariant measure, the following proposition characterizes an invariant measure of the conjugate system $S=h\circ T\circ h^{-1}$ via a pushforward measure. For completeness, we provide a short proof of this fact.
Proposition 2.1 (Invariant measures under conjugacy).
Let $X$ and $Y$ be Polish spaces, let $T:X\to X$ be continuous, and let $\mu\in\mathcal{P}(X)$ be $T$-invariant. If $h:X\to Y$ is a homeomorphism, then $h\#\mu\in\mathcal{P}(Y)$ is $S$-invariant, where $S=h\circ T\circ h^{-1}$. Moreover, $\mu$ is $T$-ergodic, then $h\#\mu$ is $S$-ergodic.
Proof.
First, note that for any Borel subset $B\subseteq Y$ it holds that
$(h\#\mu)(S^{-1}(B))=\mu(h^{-1}(S^{-1}(B)))$ $=\mu(h^{-1}(h(T^{-1}(h^{-1}(B)))))$
$=\mu(T^{-1}(h^{-1}(B)))=\mu(h^{-1}(B))=(h\#\mu)(B),$
and thus $h\#\mu$ is $S$-invariant. Furthermore, if $S^{-1}(B)=B$, then it holds that $h(T^{-1}(h^{-1}(B)))=B,$ which means that $T^{-1}(h^{-1}(B))=h^{-1}(B).$ If $\mu$ is ergodic, this directly implies that $\mu(h^{-1}(B))\in\{0,1\}$, which gives us that $(h\#\mu)(B)\in\{0,1\}$, completing the proof. ∎
Remark 2.1.
In situations when $h:X\to Y$ is a homeomorphism onto its image, it is understood that the inverse map $h^{-1}$ is only well-defined on the set $h(X)\subseteq Y$. Moreover, if $\mu$ is $T$-invariant, then $h\#\mu\in\mathcal{P}(Y)$ can still be viewed as an invariant measure of the corresponding conjugate system $S:h(X)\to h(X)$, given by $S=h\circ T\circ h^{-1}$, in the sense that its restriction to the set $h(X)\supseteq\text{supp}(h\#\mu)$ is $S$-invariant.
Proposition 2.1 motivates a key insight surrounding the issue of uniqueness when using invariant measures to compare dynamical systems. If $\mu\in\mathcal{P}(X)$ is invariant under both $T:X\to X$ and $S:X\to X$, then we cannot distinguish between the dynamical systems by studying the invariant measure alone. To resolve this challenge, we will instead make a change of coordinates and study the invariant measures of related, conjugate dynamical systems. That is, we will construct maps $h_{T},h_{S}:X\to Y$, which are homeomorphic onto their image, and consider the invariant measures, $h_{T}\#\mu\in\mathcal{P}(Y)$ and $h_{S}\#\mu\in\mathcal{P}(Y)$, of the resulting conjugate systems, $\hat{T}=h_{T}\circ T\circ h_{T}^{-1}$ and $\hat{S}=h_{S}\circ S\circ h_{S}^{-1}$. Our ability to distinguish between $T$ and $S$ via $h_{T}\#\mu$ and $h_{S}\#\mu$ depends on how the conjugating maps $h_{T}$ and $h_{S}$ are chosen. In what follows, we will argue that a powerful choice for our purpose is given by the time-delay map, originating from Takens’ seminal embedding theory [60].
The technique of time-delay embedding has revolutionized the way in which nonlinear trajectory data is studied across a diverse range of applications, including fluid mechanics [68], neuroscience [59], financial market analysis [52], and time-series prediction [66], to name a few. Rather than observing trajectories $\{T^{k}(x)\}_{k=0}^{N}$ of the full state of a dynamical system, practitioners often only have access to time-series projections of the form $\{y(T^{k}(x))\}_{k=0}^{N}$, for some scalar-valued observable $y.$ In certain situations, the method of time-delays can provide a reconstruction of the original dynamical system, up to topological conjugacy, using only this scalar time-series data.
Definition 2.6 (Time-delay map).
Consider a Polish space $X$, a map $T:X\to X$, an observable function $y:X\to\mathbb{R}$, and the time delay parameter $m\in\mathbb{N}$. The time-delay map is defined as
$\Psi_{(y,T)}^{(m)}(x):=(y(x),y(T(x)),\dots,y(T^{m-1}(x)))\in\mathbb{R}^{m},$ (2)
for each $x\in X$.
In (2), we stress the dependence of the time-delay map $\Psi_{(y,T)}^{(m)}$ on the scalar observation function $y:X\to\mathbb{R}$, the underlying system $T:X\to X$, and the dimension $m\in\mathbb{N}$. When $T$ is continuous and the time-delay map $\Psi_{(y,T)}^{(m)}$ is injective, one can build a dynamical system in the reconstruction space $\mathbb{R}^{m}$ based on the following definition.
Definition 2.7 (Delay-coordinate dynamics).
Consider a Polish space $X$, a map $T:X\to X$, an observable $y:X\to\mathbb{R}$, and the time delay parameter $m\in\mathbb{N}$. Assume that $\Psi_{(y,T)}^{(m)}:X\to\mathbb{R}^{m}$, given in (2), is injective. Then, the delay-coordinate dynamics are given by
$\widehat{T}_{(y,m)}:\Psi_{(y,T)}^{(m)}(X)\to\Psi_{(y,T)}^{(m)}(X),\qquad\widehat{T}_{(y,m)}:=\Big{[}\Psi_{(y,T)}^{(m)}\Big{]}\circ T\circ\Big{[}\Psi_{(y,T)}^{(m)}\Big{]}^{-1}.$
We remark that many approaches for numerically determining a suitable embedding dimension from time-series data, such that the delay-coordinate dynamics are well-defined, have been explored [10, 13, 36, 49, 61]. Since the delay-coordinate map $\widehat{T}_{(y,m)}$ is conjugate to the state-coordinate map $T:X\to X$, we now revisit Proposition 2.1 to motivate our definition for an invariant measure in time-delay coordinates. When $T\#\mu=\mu$, the invariant measure in time-delay coordinates should be viewed as the corresponding invariant measure of the conjugate system $\widehat{T}_{(y,m)}$ given by the pushforward of $\mu$ under the time-delay map (2.6); see Definition 2.8 below. Moreover, an illustration showing the difference between the state-coordinate invariant measure and the delay-coordinate invariant measure for the Lorenz-63 system is shown in Figure 3.
Definition 2.8 (Invariant Measure in Time-Delay Coordinates).
Assume that the delay map $\Psi_{(y,T)}^{(m)}:X\to\mathbb{R}^{m}$ in Definition 2.6 is injective. Then, the probability measure
$\hat{\mu}_{(y,T)}^{(m)}:=\Psi_{(y,T)}^{(m)}\#\mu\in\mathcal{P}(\mathbb{R}^{m})$ (3)
is the corresponding invariant measure in the time-delay coordinates. We also refer to (3) as a delay-coordinate invariant measure.
Figure 3: Visualization of the difference between state-coordinate and delay-coordinate invariant measures for the Lorenz-63 system. The top row shows a trajectory in the original state-coordinates, as well as its resulting invariant measure. The time-delay map (see Definition 2.6) can be applied to the state-coordinate trajectory to obtain the resulting trajectory in time-delay coordinates, and its pushforward can be applied to the state coordinate invariant measure to obtain the delay-coordinate invariant measure (see Definition 2.8), both of which are depicted in the bottom row.
2.3 Embedding Theory
This section contains the complete statements of two embedding theorems crucial for this work. In particular, we will study generalizations of the celebrated embedding theorems due to Whitney and Takens [60, 63]. Whitney showed that a generic smooth map from a smooth $n$-dimensional manifold into $\mathbb{R}^{2n+1}$ is an embedding, while Takens showed that the time-delay map from Definition 2.6 is generically an embedding into $\mathbb{R}^{2n+1}$ when $X=M$ is a smooth compact $n$-dimensional manifold, $T\in C^{2}(M,M)$ is a diffeomorphism, and $y\in C^{2}(M,\mathbb{R})$.
In these statements, the phrase “generic” indicates that the embedding property holds for an open and dense set of functions. Given that practitioners often have little control over the observable $y$, an important part of the delay map (2), the embedding property should hold for a large enough class of $y\in C^{2}(M,\mathbb{R})$ such that a “randomly chosen” observation function can be used to embed the dynamics with high probability.
Since an open and dense set in a topological space can have vanishingly small Lebesgue measure, the original statements by Takens and Whitney are insufficient in this regard [28, 53]. In particular, one would like to say that the time-delay map constitutes an embedding for “almost all” observation functions $y\in C^{2}(M,\mathbb{R})$. Along these lines, [53] proves generalization of both Whitney’s and Takens’ theorems using a measure-theoretic notion of genericity, known as prevalence, which is stronger than the topological genericity that appeared in the original statements [60]. We will thus develop our main results based on prevalence.
In Section 2.3.1, we provide an overview of the notion of prevalence, as well as some foundational results we will employ in this work. In Section 2.3.2, we then provide complete statements of the generalized versions of Takens’ and Whitneys’ embedding theorems due to Sauer, Yorke, and Casdagli [53]. Finally, in Section 2.3.3, we comment on the Borel measurability of the set of embeddings, a technical detail needed to prove our main results.
2.3.1 Prevalence
The definition of prevalence generalizes the notion of “Lebesgue almost everywhere” from finite-dimensional vector spaces to infinite dimensional function vector spaces [28, 43, 53]. Towards this, we first define the Lebesgue measure on an arbitrary finite-dimensional vector space. While such a definition will necessarily depend on a choice of basis for the target space, we will only be interested in whether a given set has zero or positive Lebesgue measure, and this is a property independent of one’s choice of basis [28, 56].
Definition 2.9 (Full Lebesgue measure on finite-dimensional vector spaces).
Let $E$ be a $k$-dimensional vector space and consider a basis $\{v_{1},\ldots,v_{k}\}\subseteq E$. A set $F\subseteq E$ has full Lebesgue measure if the coefficients $\big{\{}(a_{1},\dots,a_{k})\in\mathbb{R}^{k}:\sum_{i=1}^{k}a_{i}v_{i}\in F\big{\}}$ of its basis expansion have full Lebesgue measure in $\mathbb{R}^{k}$.
We will next define prevalence, which generalizes Definition 2.9 to infinite dimensional vector spaces. In particular, we define prevalence for Borel subsets of completely metrizable topological vector spaces. A topological vector space is a vector space in which the operations of vector addition and scalar multiplication are continuous functions. Moreover, the assumption of complete metrizability means that one can endow the topological vector space with a metric, for which every Cauchy sequence is convergent to a limit inside of the vector space. Additionally, this metric can always be chosen to be translation invariant.
Definition 2.10 (Prevalence).
Let $V$ be a completely metrizable topological vector space. A Borel subset $S\subseteq V$ is said to be prevalent if there is a finite-dimensional subspace $E\subseteq V$, known as a probe space, such that for each $v\in V$, it holds that $v+e\in S$ for Lebesgue almost every $e\in E$.
Intuitively, prevalence means that almost all perturbations of an element $v\in V$ by elements of a probe space $E$ necessarily belong to the prevalent set $S$. Note that Definition 2.10 reduces to Definition 2.9 if $V$ is a finite dimensional vector space. Moreover, if $V=C^{k}$, the space of functions whose $k$-th order derivative exists and is continuous, then it can be shown that a prevalent set is also dense in the $C^{k}$-topology [28]. It also holds that the finite intersection of prevalent sets remains prevalent. The proof of this fact leverages the result that any finite-dimensional subspace containing a probe space is itself a probe space. These results play crucial roles in our main theorems, so we prove them in Lemmas 2.1 and 2.2 for completeness.
Lemma 2.1.
Let $V$ be a completely metrizable topological vector space, assume $S\subseteq V$ is prevalent, and assume that $E\subseteq V$ is a probe space. If $E^{\prime}\subseteq V$ is a finite-dimensional vector space with $E\subseteq E^{\prime}$, then $E^{\prime}$ is also a probe space.
Proof.
We will denote $\text{dim}(E)=k$ and $\text{dim}(E^{\prime})=\ell\geq k$. Let $e_{1},\dots,e_{\ell}\in E^{\prime}$ be basis vectors for $E^{\prime}$, where $e_{1},\dots,e_{k}\in E$ are basis vectors for $E$. By the assumption that $E$ is a probe space, we know that for all $v\in V$ it holds that $v+\sum_{i=1}^{k}a_{i}e_{i}\in S$ for Lebesgue almost all $(a_{1},\dots,a_{k})\in\mathbb{R}^{k}$. Now, let $v\in V$ be fixed and notice that
$v+\sum_{i=1}^{\ell}a_{i}e_{i}=\Bigg{(}v+\sum_{i=k+1}^{\ell}a_{i}e_{i}\Bigg{)}+\sum_{i=1}^{k}a_{i}e_{i}.$ (4)
Due to the decomposition (4) and the fact that $E$ is a probe space, it holds for fixed $v\in V$ and $a_{k+1},\dots,a_{\ell}\in\mathbb{R}$ that $v+\sum_{i=1}^{\ell}a_{i}e_{i}\in S$ for $\lambda_{k}$-almost every $(a_{1},\dots,a_{k})\in\mathbb{R}^{k}$.
To conclude the proof, it remains to show that $v+\sum_{i=1}^{\ell}a_{i}e_{i}\in S$ for $\lambda_{\ell}$-almost every $(a_{1},\dots,a_{\ell})\in\mathbb{R}^{\ell}$, which will be a consequence of the Tonelli’s theorem [57, Theorem 3.2]. Towards this, we define the set $B:=\{(a_{1},\dots,a_{\ell}):v+\sum_{i=1}^{\ell}a_{i}e_{i}\notin S\}\subseteq\mathbb{R}^{\ell}$, and will show $\lambda_{\ell}(B)=0.$ To apply Tonelli’s theorem to $\chi_{B}$, the characteristic function of the set $B$, we must first justify that $B$ is Borel measurable. Towards this, define $f(a_{1},\dots,a_{\ell}):=v+\sum_{i=1}^{\ell}a_{i}e_{i}$, a continuous function from $\mathbb{R}^{\ell}$ to $V$. Since $S\subseteq V$ is Borel, it holds that $B=f^{-1}(S)$ is Borel as well.
For notational convenience, we will now rewrite $\chi_{B}$ as a function $\phi:\mathbb{R}^{k}\times\mathbb{R}^{\ell-k}\to\mathbb{R}$, given by
$\phi(x,y):=\chi_{B}(x_{1},\dots,x_{k},y_{1},\dots,y_{\ell-k}),\qquad x\in\mathbb{R}^{k},\qquad y\in\mathbb{R}^{\ell-k}.$
Note that for every fixed $y\in\mathbb{R}^{\ell-k}$, $\phi(x,y)=0$ for $\lambda_{k}$-almost every $x\in\mathbb{R}^{k}$. Then, it follows by Tonelli’s theorem that
[…]
which concludes the proof. ∎
It is now a direct consequence of Lemma 2.1 that the finite intersection of prevalent sets is prevalent.
Lemma 2.2.
Let $S_{1}$ and $S_{2}$ be two prevalent subsets of a completely metrizable topological vector space $V$. Then, the intersection $S_{1}\cap S_{2}$ is also prevalent.
Proof.
Let $E_{1}=\text{span}\{v_{1},\dots,v_{k}\}$ be a probe space for $S_{1}$ and $E_{2}=\text{span}\{w_{1},\dots,w_{\ell}\}$ be a probe space for $S_{2}$. Define $E^{\prime}:=\text{span}(\{v_{1},\dots,v_{k},w_{1},\dots,w_{\ell}\})$, and consequently $E_{1},E_{2}\subseteq E^{\prime}.$ Thus, by Lemma 2.1, $E^{\prime}$ is a probe space for both $S_{1}$ and $S_{2}$. This implies that $S_{1}\cap S_{2}$ is prevalent. ∎
2.3.2 Generalized Whitney’s and Takens’ Embedding Theorems
In addition to using the theory of prevalence, the results in [53] generalize the classical Whitney and Takens theorems in another crucial way. It is often the case that dynamical trajectories are asymptotic to a compact attracting set $A$, which has a fractal structure and is not a manifold. The classical Takens and Whitney theorems require that such an attractor is contained within a smooth, compact manifold of dimension $d$ to embed the dynamics in $12d+$ 1-dimensional reconstruction space. The fractal dimension $d_{A}$ of the set $A$ might be much less than the manifold dimension, i.e., $d_{A}\ll d$. In such cases, it is desirable to consider more efficient approaches that can guarantee a system reconstruction in $12d_{A}+$ 1-dimensions. Towards this, we now recall the following definition of box-counting dimension, which will serve as our notion for the dimension of a fractal set.
Definition 2.11 (Box counting dimension).
Let $A\subseteq\mathbb{R}^{n}$ be a compact set. Its box counting dimension is
$\text{boxdim}(A):=\displaystyle\lim_{\varepsilon\to 0}\frac{\log N(\varepsilon)}{\log(1/\varepsilon)},$
where $N(\varepsilon)$ is the number of boxes with side-length $\varepsilon$ required to cover $A$. When the limit does not exist, one can define the upper and lower box-counting dimensions by replacing the limit with liminf and limsup, respectively.
The following generalization of Whitney’s embedding theorem comes from [53, Theorem 2.3].
Theorem 2.1 (Fractal Whitney Embedding).
Let $A\subseteq\mathbb{R}^{n}$ be compact, $d:=\textup{boxdim}(A)$, and $m>2d$ be an integer. Then, almost every $F\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{m})$ is injective on $A$.
Remark 2.2.
The probe space for the prevalent set appearing in Theorem 2.1 can be chosen as the $nm$-dimensional space of linear maps between $\mathbb{R}^{n}$ and $\mathbb{R}^{m}.$
There are certain assumptions required on the periodic points of a dynamical system for Takens’ theorem to hold. Assumption 2.1 makes these technical assumptions concise and easy to reference in our main results. In what follows, $DT^{p}$ denotes the derivative of the $p$-fold composition map $T^{p}$.
Assumption 2.1 (Technical assumption on periodic points).
Let $T:U\to U$ be a diffeomorphism of an open set $U\subseteq\mathbb{R}^{n}$, $A\subseteq U$ be compact, and $m\in\mathbb{N}$. For each $p\leq m$:
  1. 1.
    The set $A_{p}\subseteq A$ of $p$-periodic points satisfies $\textup{boxdim}(A_{p})<p/2$.
  2. 2.
    The linearization $DT^{p}$ of each periodic orbit has distinct eigenvalues.
We remark that when the diffeomorphism $T$ is given by the time-$\tau$ flow map of a Lipschitz continuous vector field, one can choose $\tau$ sufficiently small such that Assumption 2.1 is satisfied; see [53]. The following generalization of Takens’ theorem can be found in [53, Theorem 2.7].
Theorem 2.2 (Fractal Takens’ Embedding).
Let $T:U\to U$ be a diffeomorphism of an open set $U\subseteq\mathbb{R}^{n}$, $A\subseteq U$ be compact and $m>2d$ where $d=\textup{boxdim}(A)$. Assume that the periodic points of $T$ with degree at most $m$ satisfy Assumption 2.1 on $A$. Then, it holds that $\Psi_{(y,T)}^{(m)}$ is injective on $A$, for almost all $y\in C^{1}(U,\mathbb{R}).$
Remark 2.3.
The probe space for the prevalent set appearing in Theorem 2.2 can be chosen as the space of polynomials for $n$ variables of degree at most $2m$.
We note that when the box-counting dimension does not exist, the generalized version of the Takens’ and Whitney’s embedding theorems (see Theorems 2.1 and 2.2) hold in terms of the lower box-counting dimension (see Definition 2.11). Moreover, the statements of Theorems 2.1 and 2.2 in [53] indicate that $F$ and $\Psi_{(y,T)}^{(m)}$ are also immersions on each compact subset of a smooth manifold contained in $A$, in addition to being injections. We do not need this property for our proofs, so we have left it out of the statements of Theorems 2.1 and 2.2 for simplicity.
2.3.3 On the Borel Measurability of the Set of Embeddings
It is important to understand how Theorems 2.1 and 2.2 can be precisely understood within the theory of prevalence. In particular, Definition 2.10 requires that a prevalent set is a Borel subset of completely metrizable topological vector space. While [53] does not discuss the measurability of the sets of functions appearing in Theorems 2.1 and 2.2, this property is essential for the proofs of our main results. We leverage Lemma 2.2 to construct new prevalent sets via the intersection operation. This construction is made possible by Tonelli’s theorem, which we can apply when the prevalent sets in question are Borel measurable (see our proof of Lemma 2.2).
In this section, we comment on the structure of $C^{1}(U,\mathbb{R}^{k})$ as a completely metrizable topological vector space, and we explicitly write down an appropriate translation-invariant metric on $C^{1}(U,\mathbb{R}^{k})$. We will then move on to verify that the sets of functions appearing in the statements of Theorems 2.1 and 2.2, for which the conclusions of injectivity are satisfied, are indeed Borel sets in this topology.
Remark 2.4.
For any open subset $U\subseteq\mathbb{R}^{n}$, we remark that $C^{r}(U,\mathbb{R}^{k})$ is a completely metrizable topological vector space. The construction we provide here follows [2]. In particular, consider a sequence of compact sets $K_{j}\subseteq U$, such that $K_{j}\subseteq\text{int}(K_{j+1})$ and $\bigcup_{j=1}^{\infty}K_{j}=U$. Then define the metric $\textrm{d}_{C^{r}}:C^{r}(U,\mathbb{R}^{k})\times C^{r}(U,\mathbb{R}^{k})\to[0,\infty)$ by setting
$\textrm{d}_{C^{r}}(f,g):=\sum_{j=1}^{\infty}\frac{1}{2^{j}}\frac{\textrm{d}_{j}(f,g)}{1+\textrm{d}_{j}(f,g)},\qquad\textrm{d}_{j}(f,g):=\max_{|\alpha|\leq r}\sup_{x\in K_{j}}\|\partial^{\alpha}f(x)-\partial^{\alpha}g(x)\|_{2},\qquad f,g\in C^{r}(U,\mathbb{R}^{k}),$
where $\|\cdot\|_{2}$ denotes the Euclidean $2$-norm and $\alpha\in\mathbb{N}^{n}$ is a multi-index. Then, $C^{r}(U,\mathbb{R}^{k})$ is complete with respect to the metric $\textrm{d}_{C^{r}},$ which is translation-invariant and continuous with respect to both scalar multiplication and vector addition [2, Proposition 1.25].
We will next verify that the set of smooth maps that are injective on a compact set $A\subseteq\mathbb{R}^{n}$ is indeed a Borel set with respect to the metric-topology induced by $\textrm{d}_{C^{1}}$; see Remark 2.4 .
Proposition 2.2.
Let $U\subseteq\mathbb{R}^{n}$ be an open set, $A\subseteq U$ be compact, and $k\in\mathbb{N}$. Then, the set $\{f\in C^{1}(U,\mathbb{R}^{k}):f\textup{ is injective on }A\}$ is a Borel subset of $C^{1}(U,\mathbb{R}^{k})$.
Proof.
For $\varepsilon>0$, define the set $D_{\varepsilon}:=\{(x,y)\in U\times U:\|x-y\|_{2}<\varepsilon\}$ and note that $D_{\varepsilon}$ is open. To see this, observe that $d:U\times U\to\mathbb{R}$, given by $d(x,y):=\|x-y\|_{2}$ is continuous and that $D_{\varepsilon}=d^{-1}((-\infty,\varepsilon))$ is the preimage of an open set under a continuous function, and is hence open. Next, consider the set
$\bm{B}_{\varepsilon}:=\{f\in C^{1}(U,\mathbb{R}^{k}):\|f(x)-f(y)\|_{2}>0,\,(x,y)\in(A\times A)\setminus D_{\varepsilon}\}.$
We claim that $\bm{B}_{\varepsilon}$ is an open set. To see this, we will fix $f\in\bm{B}_{\varepsilon}$ and produce an open ball of radius $\gamma>0$, which is centered at $f$ and fully contained within $\bm{B}_{\varepsilon}$. Towards this, let $K_{j}\subseteq U$ be a sequence of compact sets with $K_{j}\subseteq\text{int}(K_{j+1})$ and $\bigcup_{j\in\mathbb{N}}K_{j}=U$. Now, let $j\in\mathbb{N}$ be chosen such that $K_{j}\subseteq U$ is a compact set satisfying $A\subseteq K_{j}$. Furthermore, note that since the map $(x,y)\mapsto\|f(x)-f(y)\|_{2}$ is continuous over the compact set $(A\times A)\setminus D_{\varepsilon}$, there must exist some $a\in(0,1)$, such that $\|f(x)-f(y)\|_{2}>a>0,$ for all $(x,y)\in(A\times A)\setminus D_{\varepsilon}$. Now, choose $\gamma:=a/2^{j+1}>0$. Then, for any $g\in C^{1}(U,\mathbb{R}^{k})$ satisfying $\textrm{d}_{C^{1}}(f,g)<\gamma$, following the definition of $\textrm{d}_{C^{1}}$ (see Remark 2.4), we have that for any $x\in A$ that
$\frac{1}{2^{j}}\frac{\|f(x)-g(x)\|_{2}}{1+\|f(x)-g(x)\|_{2}}\leq\mathrm{d}_{C^{1}}(f,g)<\gamma,$
which rearranges to yield
$\|f(x)-g(x)\|_{2}<\frac{\gamma 2^{j}}{1-\gamma 2^{j}}=\frac{a/2}{1-a/2}<\frac{a}{2}.$
Thus, for all $(x,y)\in(A\times A)\setminus D_{\varepsilon}$, it follows that
$a$ $<\|f(x)-f(y)\|_{2}$
$\leq\|f(x)-g(x)\|_{2}+\|f(y)-g(y)\|_{2}+\|g(x)-g(y)\|_{2}$
$<\frac{a}{2}+\frac{a}{2}+\|g(x)-g(y)\|_{2},$
which implies $\|g(x)-g(y)\|_{2}>0.$ By definition, we have $g\in\bm{B}_{\varepsilon}$. Since $g$ was arbitrary, we have $\{g\in C^{1}(U,\mathbb{R}^{k}):\textrm{d}_{C^{1}}(f,g)<\gamma\}\subset\bm{B}_{\varepsilon}$. Since $f\in\mathbf{B}_{\varepsilon}$ was arbitrary, we verified that $\bm{B}_{\varepsilon}$ is open.
Next, we claim that
$\{f\in C^{1}(U,\mathbb{R}^{k}):f\text{ is injective on }A\}=\bigcap_{n\in\mathbb{N}}\bm{B}_{1/n}.$ (5)
If we can establish (5), then we have written the set of injective functions as a countable intersection of open sets, which will conclude our proof. We will verify that (5) holds by showing both set inclusions.
We first prove the $\subseteq$” direction. Assume that $f\in C^{1}(U,\mathbb{R}^{k})$ is injective on $A$. Defining $D:=\{(x,x):x\in A\}$, it holds by definition of injectivity that $\|f(x)-f(y)\|_{2}>0$, for all $(x,y)\in(A\times A)\setminus D$. In particular, since $D\subseteq D_{\varepsilon}$ for all $\varepsilon>0$ it holds that $\|f(x)-f(y)\|_{2}>0$, for all $(x,y)\in(A\times A)\setminus D_{\varepsilon}$, for any $\varepsilon>0$, and thus, $f\in\bm{B}_{1/n}$, for all $n\in\mathbb{N}$, i.e., $f\in\bigcap_{n\in\mathbb{N}}\bm{B}_{1/n}.$
We now prove the $\supseteq$” direction. Assume that $f\in C^{1}(U,\mathbb{R}^{k})$ satisfies $\|f(x)-f(y)\|_{2}>0$ for all $(x,y)\in(A\times A)\setminus D_{1/n}$, for each $n\in\mathbb{N}$. Let $(x_{0},y_{0})\in A\times A$ be fixed and $x_{0}\neq y_{0}$. If we choose $N_{0}\in\mathbb{N}$ to satisfy $N_{0}>1/\|x_{0}-y_{0}\|_{2}$, then it holds that $(x_{0},y_{0})\in(A\times A)\setminus D_{1/N_{0}}$ and therefore $\|f(x_{0})-f(y_{0})\|_{2}>0$, which means $f(x_{0})\neq f(y_{0}).$ Since $(x_{0},y_{0})\in A\times A$ where $x_{0}\neq y_{0}$ was chosen arbitrarily, we have by definition that $f$ is injective on $A$, which concludes the proof. ∎
We next verify that the set of observation functions $y\in C^{1}(U,\mathbb{R})$, such that the delay map $\Psi_{(y,T)}^{(k)}$ is injective on a compact set $A\subseteq U$ is also Borel in the metric-topology induced by $\textrm{d}_{C^{1}}.$ Throughout our proof, we use the fact that the convergence $\lim_{n\to\infty}\textrm{d}_{C^{r}}(f_{n},g)=0$ is equivalent to the convergence $\lim_{n\to\infty}\textrm{d}_{j}(f_{n},g)=0$, for all $j\in\mathbb{N}$; see [2, Proposition 1.23].
Proposition 2.3.
Let $U\subseteq\mathbb{R}^{n}$ be open, $A\subseteq U$ be compact, $T\in C^{1}(U,U)$ be a diffeomorphism, and $k\in\mathbb{N}$. Then, the set $\{y\in C^{1}(U,\mathbb{R}):\Psi_{(y,T)}^{(k)}\text{ is injective on }A\}$ is Borel.
Proof.
Define the operator $\bm{\Psi}:C^{1}(U,\mathbb{R})\to C^{1}(U,\mathbb{R}^{k})$ as $\bm{\Psi}(y)=\Psi_{(y,T)}^{(k)}=(y,y\circ T,\dots,y\circ T^{k-1})$ and observe that
$\{y\in C^{1}(U,\mathbb{R}):\Psi_{(y,T)}^{(k)}\text{ is injective on }A\}=\bm{\Psi}^{-1}(\{F\in C^{1}(U,\mathbb{R}^{k}):F\text{ is injective on }A\}).$
Since $\{F\in C^{1}(U,\mathbb{R}^{k}):F\text{ is injective on }A\}$ is Borel by Proposition 2.2, it suffices to show that $\bm{\Psi}$ is continuous, hence Borel measurable. Towards this, let us assume that $\lim_{\ell\to\infty}\textrm{d}_{C^{1}}(y_{\ell},y)=0$, which equivalently means $\lim_{\ell\to\infty}\textrm{d}_{j}(y_{\ell},y)=0$, for all $j\in\mathbb{N}$. Recall $\textrm{d}_{j}$ is defined in Remark 2.4, based upon a sequence of compact sets $K_{j}\subseteq U$, such that $K_{j}\subseteq\text{int}(K_{j+1})$ and $\bigcup_{j\in\mathbb{N}}K_{j}=U$. Our goal is to show that $\lim_{\ell\to\infty}\textrm{d}_{C^{1}}(\bm{\Psi}(y_{\ell}),\bm{\Psi}(y))=0$, as well.
First, let the multi-index $\alpha\in\mathbb{N}^{n}$ satisfy $|\alpha|=0$, so $\partial^{\alpha}\bm{\Psi}(y)=\bm{\Psi}(y)$. We fix $j\in\mathbb{N}$, and for each $10\leq i\leq k-$ 1 write $C_{i}:=T^{i}(K_{j})$, which is compact. Next, we select $j^{\prime}\in\mathbb{N}$, such that $C_{i}\subseteq K_{j^{\prime}}$, for all $10\leq i\leq k-$ 1. Then, it holds that
$\sup_{x\in K_{j}}\|\bm{\Psi}(y_{\ell})(x)-\bm{\Psi}(y)(x)\|^{2}_{2}=\sup_{x\in K_{j}}\sum_{i=0}^{k-1}(y_{\ell}(T^{i}(x))-y(T^{i}(x)))^{2}$ $\leq\sum_{i=0}^{k-1}\sup_{x\in K_{j}}(y_{\ell}(T^{i}(x))-y(T^{i}(x)))^{2}$
$=\sum_{i=0}^{k-1}\sup_{x\in C_{i}}(y_{\ell}(x)-y(x))^{2}$
$\leq\sum_{i=0}^{k-1}\sup_{x\in K_{j}^{\prime}}(y_{\ell}(x)-y(x))^{2}$
$\leq k\textrm{d}_{j^{\prime}}(y_{\ell},y)^{2}$
${\ell\to\infty}0.$
Now, we consider the case when $\alpha\in\mathbb{N}^{n}$ satisfies $|\alpha|=1.$ In this case, we have that
$\sup_{x\in K_{j}}\|\partial^{\alpha}\bm{\Psi}(y_{\ell})(x)-\partial^{\alpha}\bm{\Psi}(y)(x)\|^{2}_{2}$ $=\sup_{x\in K_{j}}\sum_{i=0}^{k-1}\Big{(}\partial^{\alpha}y_{\ell}(T^{i}(x))-\partial^{\alpha}y(T^{i}(x))\Big{)}^{2}$
$=\sup_{x\in K_{j}}\sum_{i=0}^{k-1}\Big{(}(\nabla y_{\ell}(T^{i}(x))-\nabla y(T^{i}(x)))\cdot\partial^{\alpha}T^{i}(x)\Big{)}^{2}$
$\leq\sup_{x\in K_{j}}\sum_{i=0}^{k-1}\|\nabla y_{\ell}(T^{i}(x))-\nabla y(T^{i}(x))\|_{2}^{2}\cdot\|\partial^{\alpha}T^{i}(x)\|_{2}^{2}$
$\leq\sum_{i=0}^{k-1}\sup_{x\in K_{j^{\prime}}}\|\nabla y_{\ell}(x)-\nabla y(x)\|_{2}^{2}\cdot\sup_{x\in K_{j}}\|\partial^{\alpha}T^{i}(x)\|_{2}^{2}$
${\ell\to\infty}0,$
where we have applied the Cauchy-Schwartz inequality and again used the fact that $C_{i}\subseteq K_{j^{\prime}}$ for each $10\leq i\leq k-$ 1. Thus, we have shown that
$\lim_{\ell\to\infty}\textrm{d}_{j}(\bm{\Psi}(y_{\ell}),\bm{\Psi}(y))=\lim_{\ell\to\infty}\max_{|\alpha|\leq 1}\sup_{x\in K_{j}}\|\partial^{\alpha}\bm{\Psi}(y_{\ell})(x)-\partial^{\alpha}\bm{\Psi}(y)(x)\|_{2}=0,\qquad\forall j\in\mathbb{N}.$
Equivalently, $\textrm{d}_{C^{1}}(\bm{\Psi}(y_{\ell}),\bm{\Psi}(y))\to 0$ as $\ell\to\infty$, which completes the proof. ∎
3 Main Results
In this section we state and prove our main results. The statements of our two main theorems appear in Section 3.1, a discussion of the results follows in Section 3.2, and the complete proofs can be found in Section 3.3.
3.1 Statements
In the following, we write $d_{\mu}:=\text{boxdim}(\text{supp}(\mu))$ to denote the box-counting dimension of the support of a probability measure $\mu\in\mathcal{P}(\mathbb{R}^{k})$. Our first result, Theorem 3.1, shows that equality of invariant measures in time-delay coordinates implies topological conjugacy of the underlying dynamical systems on the supports of their respective invariant measures.
Theorem 3.1.
Let $T,S:U\to U$ be diffeomorphisms of an open set $U\subseteq\mathbb{R}^{n}$ and suppose that $\mu\in\mathcal{P}(U)$ is $T$-invariant, $\nu\in\mathcal{P}(U)$ is $S$-invariant, and $\textnormal{supp}(\mu),\textnormal{supp}(\nu)\subseteq U$ are compact. Further assume that $m>2\max\{d_{\mu},d_{\nu}\}$, and that the periodic points of $T$ and $S$ with degree at most $m$ satisfy Assumption 2.1 on $\textnormal{supp}(\mu)$ and $\textnormal{supp}(\nu)$, respectively. Then, equality $\hat{\mu}_{(y,T)}^{(m+1)}=\hat{\nu}_{(y,S)}^{(m+1)}$ of the delay-coordinate invariant measures implies topological conjugacy of the maps $T|_{\textup{supp}(\mu)}$ and $S|_{\textup{supp}(\nu)}$, for almost every $y\in C^{1}(U,\mathbb{R})$.
A complete proof of Theorem 3.1 can be found in Section 3.3.1. It is worth noting that the required embedding dimension to form the delay-coordinate invariant measures in Theorem 3.1 is one dimension greater than the embedding dimension in Theorem 2.2. Indeed, the key to our proof is the observation that a single point
[…]
in $m+1$ dimensional time-delay coordinates determines both $\Psi_{(y,T)}^{(m)}(x)$ and $\Psi_{(y,T)}^{(m)}(T(x))$. Together, these quantities account for one forward step of the dynamics in $m$ dimensional time-delay coordinates; see Definition 2.7.
Remark 3.1.
Our proof of Theorem 3.1 primarily relies on the topology of delay coordinates, and the conclusion of Theorem 3.1 still follows if one replaces the assumption $\hat{\mu}_{(y,T)}^{(m+1)}=\hat{\nu}_{(y,S)}^{(m+1)}$ with the weaker statement that $\text{supp}(\hat{\mu}_{(y,T)}^{(m+1)})=\text{supp}(\hat{\nu}_{(y,S)}^{(m+1)})$.
While Theorem 3.1 assumed access to only a single observable $y\in C^{1}(U,\mathbb{R})$, our next main result, Theorem 3.2, shows that the invariant measures in time-delay coordinates derived from a finite set of such observables, along with a suitable initial condition, uniquely determine the dynamical system on $\text{supp}(\mu)$.
Theorem 3.2.
Let $T,S:U\to U$ be diffeomorphisms of an open set $U\subseteq\mathbb{R}^{n}$. Suppose that $\mu\in\mathcal{P}(U)$ is both $T$-invariant and $S$-invariant, and $\textnormal{supp}(\mu)\subseteq U$ is compact. Further assume that $m>2d_{\mu}$, and that the periodic points of $T$ and $S$ with degree at most $m$ satisfy Assumption 2.1 on $\textnormal{supp}(\mu)$. If
  1. (1)
    there exists $x^{*}\in B_{\mu,T}\cap\textup{supp}(\mu)$, such that $T^{k}(x^{*})=S^{k}(x^{*})$ for $11\leq k\leq m-$ 1, and
  2. (2)
    $\hat{\mu}_{(y_{j},T)}^{(m+1)}=\hat{\mu}_{(y_{j},S)}^{(m+1)}$ for $1\leq j\leq m$, where $Y:=(y_{1},\dots,y_{m})\in C^{1}(U,\mathbb{R}^{m})$ is a vector-valued observable,
then for almost every $Y\in C^{1}(U,\mathbb{R}^{m})$, it holds that $T=S$ everywhere on $\textup{supp}(\mu)$.
The complete proof of Theorem 3.2, which can be found in Section 3.3.2, leverages Theorem 3.1 together with the generalized version of Whitney’s embedding theorem (see Theorem 2.1).
3.2 Discussion
Theorems 3.1 and 3.2 are our main results for comparing dynamical systems through delay-coordinate invariant measures, based upon uniform time-delay embeddings. We remark that our results hold in great generality, as we place no assumptions on the invariant measures under consideration. In particular, these statements still hold when the invariant measures are singular with respect to the Lebesgue measure and have fractal support, a common situation amongst attracting dynamical systems [23]. Moreover, we have formulated our results using the mathematical theory of prevalence [28], arguing that the conclusions of Theorems 3.1 and 3.2 hold for almost all observation functions; see Definition 2.10. Practitioners often have little control over the measurement device for data collection, so our proposed approach of comparing dynamical systems using invariant measures in time-delay coordinates remains broadly applicable.
Figure 4: Reconstructing the dynamics of the Lorenz-63 attractor using invariant measure-based loss functions. While the loss function $\mathcal{J}_{1}$, based only on the state-coordinate invariant measure, is insufficient for reconstructing the dynamics (bottom left), the loss function $\mathcal{J}_{2},$ which enforces equality of the delay-coordinate invariant measures, yields an accurate reconstruction (bottom right). We train two models with identical initializations based upon finite-sample approximations of $\mathcal{J}_{1}$ and $\mathcal{J}_{2}$, consisting of $N=2000$ random samples. We use a time-delay of $\tau=0.01$ and an embedding dimension of $d=6$ to approximate the delay-coordinate invariant measures.
In Theorem 3.1, we show that the equality of two invariant measures in time-delay coordinates implies a topological conjugacy of the underlying dynamical systems. This represents a significant advancement compared to standard state coordinate invariant measure matching. One of the primary shortcomings of [8, 25, 65], which performs system identification by comparing state coordinate invariant measures, is the inability of such approaches to enforce ergodicity of the reconstructed system. Since ergodicity is preserved under topological conjugacy (see Proposition 2.1), the approach of comparing invariant measures in time-delay coordinates, inspired by Theorem 3.1, resolves this challenge.
We present numerical evidence highlighting this perspective in Figure 4, where we compare the ability of loss functions based on (i) the state-coordinate invariant measure and (ii) the delay-coordinate measure to reconstruct the full dynamics of the Lorenz-63 system. In particular, we parameterize $T_{\theta}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ as a fully-connected neural network, where $\theta\in\mathbb{R}^{p}$ comprises the network’s weights and biases, and we attempt to recover the dynamics of the Lorenz-63 system using the loss functions
$\mathcal{J}_{1}(\theta):=\mathcal{D}(T_{\theta}\#\mu^{*},T^{*}\#\mu^{*}),\qquad\mathcal{J}_{2}(\theta):=\mathcal{D}(T_{\theta}\#\mu^{*},T^{*}\#\mu^{*})+\mathcal{D}(\Psi_{\theta}\#\mu^{*},\Psi^{*}\#\mu^{*}),$ (7)
where $\mu^{*}\in\mathcal{P}(\mathbb{R}^{3})$ is the Lorenz-system’s invariant measure, and $\mathcal{D}:\mathcal{P}(\mathbb{R}^{3})\times\mathbb{R}^{3}\to[0,\infty)$ is a metric or divergence on the space of probability measures, which we choose as the energy distance Maximum-Mean Discrepancy [24, 18]. Moreover, in (7) the map $T^{*}:\mathbb{R}^{3}\to\mathbb{R}^{3}$ represents the ground truth time-$\tau$ flow map of the Lorenz-63 system, and $\Psi^{*}$ is the ground truth time-delay map for $T^{*}$, based on the partially observed first component of the dynamics. The map $\Psi_{\theta}$ is the delay-coordinate map parameterized by the learned dynamics $T_{\theta}$; see Definition 2.6. Notably, if $\mathcal{J}_{1}(\theta)$ is reduced to zero, then $T_{\theta}$ admits $\mu^{*}$ as an invariant measure, and if $\mathcal{J}_{2}(\theta)$ is reduced to zero, then the delay-coordinate invariant measures for $T_{\theta}$ and $T^{*}$ additionally match. The results of Figure 4 indicate that matching state-coordinate invariant measures is insufficient for reconstructing the Lorenz-63 dynamics, whereas matching the delay-coordinate invariant measures can yield a successful reconstruction.
It is also worth noting that the invariant measures in delay coordinates used in Theorem 3.1 can be constructed entirely through partial observations of the full state. This property is essential in large-scale applications when the full state is not directly observable. However, in certain situations when one can probe the full state, Theorem 3.2 tells us that the delay invariant measures corresponding to a finite set of observation functions contain sufficient information to reconstruct the underlying dynamics, provided that a suitable initial condition also holds. We remark that the initial condition $T^{k}(x^{*})=S^{k}(x^{*})$, for $11\leq k\leq m-$ 1, appearing in Theorem 3.2 is relatively mild, only requiring that the two systems agree for finitely many iterations at some initial condition $x^{*}\in\text{supp}(\mu)\cap B_{\mu,T}.$
Rather than considering the invariant measures in time-delay coordinates from Theorem 3.2, it may be conceptually simpler to instead study $(Y,\,Y\circ T)\#\mu\in\mathcal{P}(\mathbb{R}^{2m}),$ where $Y\in C^{1}(U,\mathbb{R}^{m})$ is a vector-valued observable. By Proposition 3.1, whose proof is analogous to Theorem 3.1 and appears in Section 3.3.3, this measure contains sufficient information to recover $T$ on $\text{supp}(\mu)$, for almost all $Y\in C^{1}(U,\mathbb{R}^{m})$.
Proposition 3.1.
Let $T,S:U\to U$ be diffeomorphisms of an open set $U\subseteq\mathbb{R}^{n}$, suppose that $\mu\in\mathcal{P}(U)$ is both $T$-invariant and $S$-invariant, and assume that $\textnormal{supp}(\mu)\subseteq U$ is compact. Then, the equality $(Y,Y\circ T)\#\mu=(Y,Y\circ S)\#\mu$ implies $T=S$ on $\textup{supp}(\mu)$, for almost all $Y\in C^{1}(U,\mathbb{R}^{m})$, where $m>2d_{\mu}$.
Theorem 3.2 and Proposition 3.1 motivate two different strategies for uniquely recovering the underlying dynamical system using measure-based comparisons. In particular, we can choose to study either
[…]
in order to uniquely recover the underlying system. Given that evaluating metrics on the space of probability measures suffers from the curse of dimensionality [44], the former approach based upon Theorem 3.2 and the delay-coordinate invariant measure is likely to be more computationally feasible.
3.3 Proof of Main Results
We are now ready to present the proofs of our main results. The proof of Theorem 3.1, which is a consequence of the generalized Takens embedding theorem (Theorem 2.2) combined with an observation about the topology of time-delay coordinate systems, appears in Section 3.3.1. In Section 3.3.2, we then leverage Theorem 3.1, the generalized Takens embedding theorem (Theorem 2.2), and the generalized Whitney embedding theorem (Theorem 2.1), to prove Theorem 3.2. The proof of Proposition 3.1 then appears in Section 3.3.3.
3.3.1 Proof of Theorem 3.1
In this section, we present a complete proof of Theorem 3.1. We begin by establishing a useful lemma that relates the support of a probability measure to the support of its pushforward under a continuous mapping.
Lemma 3.1.
Let $X$ and $Y$ be Polish spaces and $\mu\in\mathcal{P}(X)$ be a Borel probability measure with compact support, and assume that $f:X\to Y$ is continuous. Then, $f(\textup{supp}(\mu))=\textup{supp}(f\#\mu).$
Proof.
We prove the result via double inclusion, beginning first with the $\supseteq$” direction. Note that since $f$ is continuous and $\text{supp}(\mu)$ is compact in $X$, it holds that $D:=f(\text{supp}(\mu))$ is compact in $Y$. Moreover, using the fact that $\text{supp}(\mu)\subseteq f^{-1}(D)$, we obtain that
$(f\#\mu)(D)=\mu(f^{-1}(D))\geq\mu(\text{supp}(\mu))=1.$
Since $D$ is closed in $Y$, this implies that $\text{supp}(f\#\mu)\subseteq D=f(\text{supp}(\mu))$. To prove the $\subseteq$” inclusion, we note by continuity that $C:=f^{-1}(\text{supp}(f\#\mu))$ is a closed subset of $X$ which satisfies
$\mu(C)=\mu(f^{-1}(\text{supp}(f\#\mu)))=(f\#\mu)(\text{supp}(f\#\mu))=1.$
Thus, $\text{supp}(\mu)\subseteq C=f^{-1}(\text{supp}(f\#\mu)),$ which implies that
$f(\text{supp}(\mu))\subseteq f(f^{-1}(\text{supp}(f\#\mu))\subseteq\text{supp}(f\#\mu).$
This establishes the inclusion $f(\text{supp}(\mu))\subseteq\text{supp}(f\#\mu)$ and completes the proof. ∎
The conclusion of Theorem 2.2 states that the time-delay map $\Psi_{(y,T)}^{(m)}:U\to\mathbb{R}^{m}$ is injective on a compact subset $A\subseteq\mathbb{R}^{n}$. The following lemma will allow us to conclude that $\Psi_{(y,T)}^{(m)}:A\to\Psi_{(y,T)}^{(m)}(A)$ is in-fact a homeomorphism. This fact will be needed when we construct the conjugating map (see Definition 2.5) appearing in the conclusion of Theorem 3.1.
Lemma 3.2.
([58, Proposition 13.26]) Let $X$ and $Y$ be compact spaces and assume that $f:X\to Y$ is continuous and invertible. Then, $f$ is a homeomorphism, i.e., $f^{-1}$ is also continuous.
It is a well-known fact that the graph $\mathcal{G}[T]:=\{(x,T(x)):x\in X\}$ of a continuous function $T:X\to X$ uniquely characterizes $T$ at all points in $X$. Indeed, if $\mathcal{G}[T]=\mathcal{G}[S]$ for some continuous $S:X\to X$, then one can easily deduce that $T=S.$ Our proof of Theorem 3.1 can be viewed as a generalization of this observation to the case when the graph $\mathcal{G}[T]$ is distorted by the time-delay map; see Definition 2.6.
Proof of Theorem 3.1.
By Theorem 2.2, we have that
$\mathcal{Y}_{1}$ :=$:=\{y\in C^{1}(U,\mathbb{R}):\Psi_{(y,S)}^{(m)}\text{ is injective on }\text{supp}(\nu)\},$
$\mathcal{Y}_{2}$ :=$:=\{y\in C^{1}(U,\mathbb{R}):\Psi_{(y,T)}^{(m)}\text{ is injective on }\text{supp}(\mu)\},$
are prevalent subsets of $C^{1}(U,\mathbb{R})$. By Lemma 2.2, it holds that the intersection $\mathcal{Y}:=\mathcal{Y}_{1}\cap\mathcal{Y}_{2}$ is also prevalent in $C^{1}(U,\mathbb{R})$. Now, let $y\in\mathcal{Y}$ be fixed and assume that
$\Psi_{(y,S)}^{(m+1)}\#\nu=\Psi_{(y,T)}^{(m+1)}\#\mu.$ (8)
Since the mappings
$\Psi_{(y,S)}^{(m)}\Big{|}_{\text{supp}(\nu)}:\text{supp}(\nu)\to\Psi_{(y,S)}^{(m)}(\text{supp}(\nu)),\qquad\Psi_{(y,T)}^{(m)}\Big{|}_{\text{supp}(\mu)}:\text{supp}(\mu)\to\Psi_{(y,T)}^{(m)}(\text{supp}(\mu))$
are continuous, invertible, and the sets $\text{supp}(\nu)$ and $\text{supp}(\mu)$ are compact, it follows from Lemma 3.2 that the map $\Theta_{y}:\text{supp}(\nu)\to\text{supp}(\mu)$, given by
$\Theta_{y}(x):=\Bigg{(}\Big{[}\Psi_{(y,T)}^{(m)}\Big{]}\bigg{|}_{\text{supp}(\mu)}^{-1}\circ\Big{[}\Psi_{(y,S)}^{(m)}\Big{]}\bigg{|}_{\text{supp}(\nu)}\Bigg{)}(x),\qquad\forall x\in\text{supp}(\nu),$ (9)
is a well-defined homeomorphism. We now aim to show that $T|_{\text{supp}(\mu)}$ and $S|_{\text{supp}(\nu)}$ are topologically conjugate (see Definition 2.5) via the homeomorphism $\Theta_{y}$.
Returning to analyzing (8), we have by Lemma 3.1 that
$\Psi_{(y,S)}^{(m+1)}(\text{supp}(\nu))=\text{supp}\Big{(}\Psi_{(y,S)}^{(m+1)}\#\nu\Big{)}=\text{supp}\Big{(}\Psi_{(y,T)}^{(m+1)}\#\mu\Big{)}=\Psi_{(y,T)}^{(m+1)}(\text{supp}(\mu)).$ (10)
Now, let $x\in\text{supp}(\nu)$ be fixed, and note that by the equality of sets (10) and the definition of the time-delay map (see Definition 2.6), there must exist some $z\in\text{supp}(\mu)$, such that
[…]
By equating the first $m$ components, and then the last $m$ components, of the $m+1$ dimensional vectors appearing in (11), we obtain the following two equalities:
$\Psi_{(y,S)}^{(m)}(x)$ $=\Psi_{(y,T)}^{(m)}(z),$ (12)
$\Psi_{(y,S)}^{(m)}(S(x))$ $=\Psi_{(y,T)}^{(m)}(T(z)).$ (13)
Since $x\in\text{supp}(\nu)$ and $z\in\text{supp}(\mu)$, we deduce from (12) that $z=\Theta_{y}(x)$. Substituting this equality into (13) yields
$\Psi_{(y,S)}^{(m)}(S(x))=\Psi_{(y,T)}^{(m)}(T(\Theta_{y}(x)))$ (14)
Moreover, since $T\#\mu=\mu$ and $S\#\nu=\nu$, it follows by Lemma 3.1 that $T(\text{supp}(\mu))=\text{supp}(\mu)$ and $S(\text{supp}(\nu))=\text{supp}(\nu)$. Thus, $S(x)\in\text{supp}(\nu)$ and $T(\Theta_{y}(x))\in\text{supp}(\mu)$. This allows us to rearrange (14) to find
$S(x)=(\Theta_{y}^{-1}\circ T\circ\Theta_{y})(x).$
Since $x\in\text{supp}(\nu)$ was arbitrary, we have the general equality
$S|_{\text{supp}(\nu)}=\Theta_{y}^{-1}\circ T|_{\text{supp}(\mu)}\circ\Theta_{y},$
which completes the proof. ∎
3.3.2 Proof of Theorem 3.2
This section contains a proof of Theorem 3.2. We begin by establishing several lemmas which are needed in our proof of the result. First, we will consider the case when two systems agree along an orbit, i.e., $S^{k}(x^{*})=T^{k}(x^{*})$ for all $k\in\mathbb{N}$. We will show that when $x^{*}\in B_{\mu,T}$ that the equality of $S$ and $T$ along the orbit initiated at $x^{*}$ implies that $S$ and $T$ agree everywhere on $\text{supp}(\mu)$.
Lemma 3.3.
Let $S,T:U\to U$ be continuous maps on an open set $U\subseteq\mathbb{R}^{n}$, let $\mu\in\mathcal{P}(U)$, and let $\textup{supp}(\mu)\subseteq U$ be compact. If for some $x^{*}\in B_{\mu,T}$ it holds that $S^{k}(x^{*})=T^{k}(x^{*})$ for all $k\in\mathbb{N}$, then $S|_{\textup{supp}(\mu)}=T|_{\textup{supp}(\mu)}$.
Proof.
Define the continuous function $\phi\in C(U)$ by setting $\phi(x):=\|S(x)-T(x)\|_{2}\geq 0$, for each $x\in U.$ Since $x^{*}\in B_{\mu,T}$, it follows from Definition 2.4 that
[…]
Moreover, since $T^{k}(x^{*})=S^{k}(x^{*})$ for all $k\in\mathbb{N}$, we find
$\lim_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}\phi(T^{k}(x^{*}))$ $=\lim_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}\|S(T^{k}(x^{*}))-T^{k+1}(x^{*})\|_{2}$
$=\lim_{N\to\infty}\frac{1}{N}\sum_{k=0}^{N-1}\|S^{k+1}(x^{*}))-T^{k+1}(x^{*})\|_{2}$
$=0.$
Since $\|S(x)-T(x)\|_{2}\geq 0$ for all $x\in U$, it must hold that $S(x)=T(x)$ for $\mu$-almost all $x\in U$.
We now define $C:=\{x\in U:S(x)=T(x)\}$ and note $C=\{x\in U:S(x)-T(x)=0\}=(S-T)^{-1}(\{0\}).$ Since $S-T$ is continuous, it holds that $C$ is closed. Moreover, since $\mu(C)=1$ it follows by Definition 2.1 that $\text{supp}(\mu)\subseteq C$. Thus, $S|_{\textup{supp}(\mu)}=T|_{\textup{supp}(\mu)},$ as wanted. ∎
In our proof of Theorem 3.2, we will require that $\Psi_{(y_{i},T)}^{(m)}$ is injective on $\text{supp}(\mu)$ for each $1\leq i\leq m$, where $Y=(y_{1},\dots,y_{m})\in C^{1}(U,\mathbb{R}^{m}).$ The following result (Lemma 3.4) is used to show that the set of all $Y\in C^{1}(U,\mathbb{R}^{m})$ for which this property holds is prevalent.
Lemma 3.4.
Assume that $\mathcal{Y}\subseteq C^{1}(U,\mathbb{R})$ is prevalent. Then, the set of functions $\bm{\mathcal{Y}}:=\{(y_{1},\dots,y_{m}):y_{i}\in\mathcal{Y}\}$ is prevalent in $C^{1}(U,\mathbb{R}^{m}).$
Proof.
We first show that $\bm{\mathcal{Y}}\subseteq C^{1}(U,\mathbb{R}^{m})$ is Borel. Towards this, define the projection
$\bm{\pi}_{i}:C^{1}(U,\mathbb{R}^{m})\to C^{1}(U,\mathbb{R}),\qquad\bm{\pi}_{i}((y_{1},\dots,y_{m})):=y_{i}\in C^{1}(U,\mathbb{R}),$
for all $(y_{1},\dots,y_{m})\in C^{1}(U,\mathbb{R}^{m})$ and each $1\leq i\leq m.$ The projection $\bm{\pi}_{i}$ is continuous, hence Borel measurable. Since $\mathcal{Y}$ is Borel, it holds that $\bm{\pi}_{i}^{-1}(\mathcal{Y})$ is also Borel for $1\leq i\leq m$. Then, we can write $\bm{\mathcal{Y}}=\bigcap_{i=1}^{m}\bm{\pi}_{i}^{-1}(\mathcal{Y})$, which verifies that $\bm{\mathcal{Y}}\subseteq C^{1}(U,\mathbb{R}^{m})$ is Borel.
Since $\mathcal{Y}$ is prevalent, there exists a $k$-dimensional probe space $E\subseteq C^{1}(U,\mathbb{R})$ admitting a basis $\{e_{i}:1\leq i\leq k\}$, such that for any $y\in C^{1}(U,\mathbb{R})$, it holds that $y+\sum_{i=1}^{k}a_{i}e_{i}\in\mathcal{Y}$ for Lebesgue almost every $(a_{1},\dots,a_{k})\in\mathbb{R}^{k}$. Next, we will define the augmented probe space
$\bm{E}:=\{(v,\dots,v):v\in E\}\subseteq C^{1}(U,\mathbb{R}^{m}),$
and note that $\{(e_{i},\dots,e_{i}):1\leq i\leq k\}$ constitutes a basis for $\bm{E}.$ Now, fix $(y_{1},\dots,y_{m})\in C^{1}(U,\mathbb{R}^{m})$ and observe that for each $1\leq j\leq m$, there exists a full Lebesgue measure set $B_{j}\subseteq\mathbb{R}^{k}$, such that $y_{j}+\sum_{i=1}^{k}a_{i}e_{i}\in\mathcal{Y}$ for all $(a_{1},\dots,a_{k})\in B_{j}$. By construction, the set $B:=\bigcap_{j=1}^{k}B_{j}$ has full Lebesgue measure in $\mathbb{R}^{k}$, and for all $(a_{1},\dots,a_{k})\in B$ it then holds that
$(y_{1},\dots,y_{m})+\sum_{i=1}^{k}a_{i}(e_{i},\dots,e_{i})=\Bigg{(}y_{1}+\sum_{i=1}^{k}a_{i}e_{i},\dots,y_{m}+\sum_{i=1}^{k}a_{i}e_{i}\Bigg{)}\in\bm{\mathcal{Y}},$
which completes the proof.
In our proof of Theorem 3.2, we will utilize the generalized Takens’ theorem (Theorem 2.2) to conclude that each $\Psi_{(y_{i},T)}^{(m)}$ is injective on the compact set $\text{supp}(\mu)$, for each $1\leq i\leq m$. We would also like to use the generalized Whitney theorem (Theorem 2.1) to conclude that $Y=(y_{1},\dots,y_{m})$ is injective on $\text{supp}(\mu)$. However, in the generalized Whitney theorem it is assumed that $Y\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{m})$, whereas in our statement of Theorem 3.2 we have $Y\in C^{1}(U,\mathbb{R}^{m})$, for an arbitrary open set $U\subseteq\mathbb{R}^{n}$. The following lemma leverages the Whitney extension theorem (see [64]) to provide a reformulation of the generalized Whitney embedding theorem in this setting.
Lemma 3.5.
Let $A\subseteq U\subseteq\mathbb{R}^{n},$ where $A$ is compact and $U$ is open, set $d:=\textnormal{boxdim}(A)$, and let $m>2d$ be an integer. For almost every smooth map $F\in C^{1}(U,\mathbb{R}^{m})$, it holds that $F$ is one-to-one on $A$ and an immersion on each compact subset of a smooth manifold contained in $A$.
Proof.
By Theorem 2.1, the result holds when $U=\mathbb{R}^{n}$, and by Remark 2.2 a suitable probe is given by space of linear maps between $\mathbb{R}^{n}$ and $\mathbb{R}^{m}$. Hereafter, we will denote this space by $\mathcal{L}(\mathbb{R}^{n},\mathbb{R}^{m})$. In the case when $U\neq\mathbb{R}^{n}$, we claim that the restricted space of linear maps $E:=\{L|_{U}:L\in\mathcal{L}(\mathbb{R}^{n},\mathbb{R}^{m})\}$ is a suitable probe. We remark that if $W:=\{L_{1},\dots,L_{nm}\}$ is a basis for $\mathcal{L}(\mathbb{R}^{n},\mathbb{R}^{m})$, which is $nm$-dimensional, then $\hat{W}:=\{L_{1}|_{U},\dots,L_{nm}|_{U}\}$ is a basis for $E$, which remains $nm$-dimensional.
To verify that $\hat{W}$ is a basis, let $\hat{L}\in E$, and note that $\hat{L}=L|_{U}$ for some $L\in\mathcal{L}(\mathbb{R}^{n},\mathbb{R}^{m})$. Since $W$ is a basis for $\mathcal{L}(\mathbb{R}^{n},\mathbb{R}^{m})$, we can write $L=\sum_{i=1}^{nm}a_{i}L_{i}$ for some coefficients $a_{i}\in\mathbb{R}$, $1\leq i\leq nm$, and thus $\hat{L}=\sum_{i=1}^{nm}a_{i}L_{i}|_{U}.$ Therefore, $\hat{W}$ spans $E$. To see that the elements of $\hat{W}$ are linearly dependent, assume that $\sum_{i=1}^{nm}a_{i}L_{i}|_{U}\equiv 0\in\mathbb{R}^{m}$ for some coefficients $a_{i}\in\mathbb{R}$ with $1\leq i\leq nm$. By linearity and the fact that $U$ is open, this implies $\sum_{i=1}^{nm}a_{i}L_{i}\equiv 0\in\mathbb{R}^{m}$, and since $\{L_{i}\}_{i=1}^{nm}$ forms a basis, this implies $a_{i}=0$ for each $1\leq i\leq nm$.
To complete the proof, it remains to show that $E$ is a probe space. Towards this, let $F\in C^{1}(U,\mathbb{R}^{m})$ be fixed, and note that by the Whitney Extension Theorem there exists $\tilde{F}\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{m})$, such that $\tilde{F}|_{A}=F|_{A}$; see [64]. Then, by Theorem 2.1 and Remark 2.2, it holds that $\tilde{F}+\sum_{i=1}^{nm}a_{i}L_{i}$ is injective on $A$ for Lebesgue almost all $(a_{1},\dots,a_{nm})\in\mathbb{R}^{nm}$. Since $\tilde{F}|_{A}=F|_{A}$, it follows that $F+\sum_{i=1}^{nm}a_{i}L_{i}|_{U}$ is injective on $A$ for Lebesgue almost all $(a_{1},\dots,a_{nm})\in\mathbb{R}^{nm}$, which completes the proof. ∎
We are now ready to present the proof of Theorem 3.2.
Proof of Theorem 3.2.
By Theorem 3.1, the set of $y\in C^{1}(U,\mathbb{R})$ such that the equality of measures $\hat{\mu}_{(y,S)}^{(m+1)}=\hat{\mu}_{(y,T)}^{(m+1)}$ implies the topological conjugacy of $T|_{\text{supp}(\mu)}$ and $S|_{\text{supp}(\mu)}$ is prevalent. We denote this set by $\mathcal{Y}$. Moreover, by Lemma 3.4 it holds that the set $\bm{\mathcal{Y}}:=\{(y_{1},\dots,y_{m}):y_{i}\in\mathcal{Y}\}$ is prevalent in $C^{1}(U,\mathbb{R}^{m}).$ By Lemma 3.5, it follows that the set $\bm{\mathcal{Z}}\subseteq C^{1}(U,\mathbb{R}^{m})$ of smooth maps that are injective on $\text{supp}(\mu)$ is also prevalent. Since the finite intersection of prevalent sets is prevalent (see Lemma 2.2) it holds that $\bm{\mathcal{W}}:=\bm{\mathcal{Y}}\cap\bm{\mathcal{Z}}$ is a prevalent subset of $C^{1}(U,\mathbb{R}^{m})$.
We now fix an element $Y=(y_{1},\dots,y_{m})\in\bm{\mathcal{W}}$. Note that by the conclusion of Theorem 3.1, we have
$S|_{\textup{supp}(\mu)}=(\Theta_{y_{i}}^{-1}\circ T|_{\text{supp}(\mu)}\circ\Theta_{y_{i}}),\qquad 1\leq i\leq m,$ (15)
where $\Theta_{y_{i}}:\text{supp}(\mu)\to\text{supp}(\mu)$ is defined in (9). Evaluating (15) at $x^{*}\in B_{\mu,T}\cap\text{supp}(\mu)$ and composing on both sides then yields
$S|_{\textup{supp}(\mu)}^{k}(x^{*})=(\Theta_{y_{i}}^{-1}\circ T|_{\text{supp}(\mu)}^{k}\circ\Theta_{y_{i}})(x^{*}),\qquad k\in\mathbb{N},\qquad 1\leq i\leq m.$
Then, since $S^{k}(x^{*})=T^{k}(x^{*})$ for $10\leq k\leq m-$ 1, it holds that $\Theta_{y_{i}}(x^{*})=x^{*}$ for $1\leq i\leq m$, which is a consequence of the definition of the delay map (2) and the construction of $\Theta_{y_{i}}$; see (9). Therefore,
$S|_{\textup{supp}(\mu)}^{k}(x^{*})=(\Theta_{y_{i}}^{-1}\circ T|_{\text{supp}(\mu)}^{k})(x^{*}),\qquad k\in\mathbb{N},\qquad 1\leq i\leq m.$ (16)
Using the definition of $\Theta_{y_{i}}$, we now rearrange (16) to find that
$\Psi_{(y_{i},S)}^{(m)}(S|_{\textup{supp}(\mu)}^{k}(x^{*}))=\Psi_{(y_{i},T)}^{(m)}(T|_{\textup{supp}(\mu)}^{k}(x^{*})),\qquad k\in\mathbb{N},\qquad 1\leq i\leq m.$ (17)
Again, using the definition of the delay map (2), and equating the first components of the vectors in (17) reveals that $y_{i}(S^{k}(x^{*}))=y_{i}(T^{k}(x^{*}))$ for all $i=1,\ldots,m$ and any $k\in\mathbb{N}.$ Recall that $Y=(y_{1},\ldots,y_{m})\in\bm{\mathcal{W}}=\bm{\mathcal{Y}}\cap\bm{\mathcal{Z}}$ is injective. Thus, $S^{k}(x^{*})=T^{k}(x^{*})$ for all $k\in\mathbb{N}$. The conclusion then follows from Lemma 3.3, which completes the proof. ∎
3.3.3 Proof of Proposition 3.1
Rather than using each observation function $y_{i}$ of $Y=(y_{1},\dots,y_{m})$ to construct a different delay-coordinate invariant measure $\hat{\mu}_{(y_{i},T)}^{(m+1)},$ one can additionally study the measure $(Y,Y\circ T)\#\mu$ to uniquely identify the dynamics $T$ on $\text{supp}(\mu).$ This is the content of Proposition 3.1, which we prove here for completeness. The argument is analogous to our proof of Theorem 3.1.
Proof of Proposition 3.1.
It follows by Lemma 3.4 that the set
$\mathcal{Y}:=\{Y\in C^{1}(U,\mathbb{R}^{m}):Y\text{ is injective on }{\text{supp}(\mu)}\}$
is prevalent in $C^{1}(U,\mathbb{R}^{m})$. Now, let $Y\in\mathcal{Y}$ be fixed and notice by Lemma 3.1 that
$(Y,Y\circ T)(\text{supp}(\mu))=\text{supp}((Y,Y\circ T)\#\mu)=\text{supp}((Y,Y\circ S)\#\mu)=(Y,Y\circ S)(\text{supp}(\mu)),$
and thus
$\{(Y(x),Y(T(x))):x\in\text{supp}(\mu)\}=\{(Y(x),Y(S(x))):x\in\text{supp}(\mu)\}.$ (18)
Thus, for fixed $x\in\text{supp}(\mu)$ it follows by (18) that there is $z\in\text{supp}(\mu)$ such that
$(Y(x),Y(T(x)))=(Y(z),Y(T(z))).$
Analyzing the first components we have $Y(x)=Y(z)$, and since $Y$ is injective on $\text{supp}(\mu)$ it holds that $x=z.$ Thus, $Y(T(x))=Y(S(x))$, which again the injectivity of $Y$ on $\text{supp}(\mu)$ implies that $T(x)=S(x)$. Since $x\in\text{supp}(\mu)$ was chosen arbitrarily, we have $T|_{\text{supp}(\mu)}=S|_{\text{supp}(\mu)}$. ∎
4 Conclusions
In recent years, measure-theoretic approaches have emerged as essential tools for modeling dynamical systems from data facing obstructions such as sparsity, noise, and uncertainty. While many researchers have leveraged the use of the state-coordinate invariant measure to perform system identification [8, 14, 25, 29, 55, 65], the invariant measure alone cannot distinguish between large classes of models which all share the same asymptotic behavior. To this end, we have developed a theoretical framework that justifies the use of invariant measures in time-delay coordinates for comparing dynamical systems. Notably, our approach retains the benefits of a measure-theoretic approach while ensuring the unique identifiability of the underlying dynamical system, up to a topological conjugacy. We expect our theoretical advances will inspire further work leveraging invariant measures in time-delay coordinates to perform more effective system identification in the face of data uncertainty and noise.
Acknowledgments
J. B.-G. was supported by a fellowship award under contract FA9550-21-F-0003 through the National Defense Science and Engineering Graduate (NDSEG) Fellowship Program, sponsored by the Air Force Research Laboratory (AFRL), the Office of Naval Research (ONR) and the Army Research Office (ARO). Y. Y. acknowledges support from National Science Foundation through grant DMS-2409855 and ONR through grant N00014-24-1-2088. R. M. acknowledges support for initial problem concept formulation under Air Force Office of Scientific Research (AFOSR) grant No. FA9550-23RQCOR007.